# SketchVLM: Vision language models can annotate images to explain thoughts and guide users

### 23 Apr 2026 v1

@@ 1 Introduction
Figure 1: For complex questions, modern chatbots like ChatGPT often return long text responses (a) that are hard for users to understand, verify, and follow. In contrast, SketchVLM guides users (b) step-by-step by annotating the input image and grounding answers to relevent image regions—here, guiding a user how to check their car’s oil level (source: https://www.youtube.com/watch?v=tNNyu9S65E4).
@@ 2 Related Work
Fine-tuned sketching models   MVoT [23] fine-tunes Chameleon [42] to generate interleaved text and image reasoning traces, visualizing intermediate states to support multi-step spatial reasoning. LatentSketchpad [53] and DeepSketcher [52] both move visual reasoning into learned latent or embedding spaces. LatentSketchpad is built with Gemma3 and Qwen2.5-VL-7B, while DeepSketcher uses Qwen2.5-VL-7B. ViLaSR [48] post-trains Qwen-2.5-VL-7B to sketch on the input image with an SVG overlay before responding. Similarly, ThinkMorph is fine-tuned from BAGEL-7B-MoT [15] to generate visual sketches that support its answers, though it directly edits the input image rather than overlaying SVG annotations. Unlike these approaches, we build a harness around SotA VLMs to enable them to annotate on input images by generating a layer of SVGs. Therefore, different from the literature, our approach is training-free and enjoy the generalizability of SotA VLMs to new domains.
@@ 4 Evaluation
7. Ball Drop [Uncaptioned image]   contains 198 synthetically generated images with harder to guess answers and ground-truth ball trajectories produced using PHYRE [4]. We generate equal numbers of images with 1, 2, and 3 randomly placed lines, and randomize the ball’s horizontal position. Models must predict the landing container out of four choices and trace the ball trajectory (Secs.˜B.6, 10, D3 and D2).
@@ 5.3 SketchVLM improves counting accuracy [Uncaptioned image]
Experiment   We compare SketchVLMs’ stroke-based annotations against the baseline models that directly outputs parameters for shape locations and shape size (Sec.˜4.1). To evaluate oval annotations, we convert the rendered shape into its tight enclosing bounding box before computing metrics. Performance is measured using Average Precision (AP) at an IoU threshold of 0.5.
Results   [Uncaptioned image] sketchVLM with stroke based rectangles is effective for medium and large objects, but remains limited for small-object detection. Sketch-based rectangles slightly improve performance on medium (+0.6) and large objects (+1.4), but significantly degrade small-object detection (-10.1), reducing overall AP50 from 63.1 to 58.8 (Figs.˜6, D10 and C4).
We observe that stroke-based outputs in SketchVLMs underperform the original model on small objects (Tab.˜C4). We examine detection statistics and prompt ablations and find that SketchVLMs match the original model in precision but exhibits lower recall (Tab.˜C5) and does not significantly change with variations in the drawing prompt (Tab.˜C6). Therefore, in order to get the accuracy of pre-defined shapes as well as the expressiveness of free-form strokes, we allow the model to output both simultaneously, such as in Fig.˜1.
@@
Results   SketchVLM improves part labeling for [Uncaptioned image] (+1.3) but slightly underperforms [Uncaptioned image] at strict boundary matching (-3.8) (Tab.˜4). For [Uncaptioned image], SketchVLM becomes increasingly robust under boundary dilation [11], achieving higher accuracy as tolerance increases, with remaining errors dominated by wrong-position mistakes (>79%; Tab.˜C7). For [Uncaptioned image], the gap narrows from (-3.8) at $r=0$ to (-0.6) at $r=7$, reaching near parity under modest tolerance (Tab.˜4). These errors largely correspond to small boundary offsets that are visually negligible (Figs.˜7, 8 and D11), where the original models produce more missing-label errors and SketchVLMs more position errors (Tab.˜C7). These findings indicate that SketchVLMs primarily introduce minor spatial shifts rather than semantic labeling failures, and remains competitive under reasonable boundary tolerance.
@@ 5.6 SketchVLMs outperform models fine-tuned directly on path-tracing tasks [Uncaptioned image]
Results   We find that [Uncaptioned image] and [Uncaptioned image] perform poorly on VPCT with accuracies of 37% and 27%, which are near the random choice baseline of 33.3%. On our Ball Drop dataset, they have accuracies of 35.9% and 30.3%, only slightly above the random choice baseline of 25.0%, showing how these models fail to generalize to tasks that they are not trained on. $\bm{+} +\bm{+} [Uncaptioned image]$ often removes ledges in the image (Fig.˜D1) and performs significantly worse than the baseline [Uncaptioned image] accuracy (Tab.˜2). In contrast, SketchVLMs output coherent annotations and have high accuracy and [Uncaptioned image] sketchVLM reaches 96.0% accuracy on VPCT and 79.7% on Ball Drop (Fig.˜10).
@@ 5.10 VLM judge ratings positively correlate with human judgments
We validate our annotation quality ratings from our VLM judge (Gemini-3-Flash-Preview [12] (Tab.˜F1)) against three human annotators on 2,250 annotations across Ball Drop, VPCT, and Maze Navigation using a quadratic Kappa score and Pearson Correlation. Agreement between humans and the VLM judge is moderate (quadratic Kappa of $0.51\pm 0.02$, Pearson correlation of $0.52\pm 0.01$) compared to high inter-human agreement (quadratic Kappa of $0.84\pm 0.04$, Pearson correlation of $0.85\pm 0.04$), further details are in Appendix˜E. While the VLM judge is not as consistent as the human evaluators, its ratings are positively correlated with human judgments across all tasks and models. We therefore treat the VLM judge scores as a useful proxy metric for quick and cost-effective evaluation that should be taken into consideration with both task accuracy and annotation–text alignment.
@@ B.5 Maze Navigation
Given a start point, an end point, and a set of direction commands (e.g., Up, Down, Left, Right), the model must determine if the path reaches the goal without crossing any border walls. We create 200 unique 3x3 grids where the shortest path length between the starting green square and ending red squares varies between 3 to 8 steps. For each maze, we take the ground truth path and randomly change one of the direction steps in order to make an invalid path (For example, Left, Right, Down could be changed to Right, Right, Down) (Fig.˜9).
@@ D.6 Part Labeling
Figure D11: Qualitative comparison on part labelling tasks between Refer to caption, Refer to caption sketchVLM , and Refer to caption.
@@ D.8 Connect-the-Dots: Bézier Curves versus Lines
Figure D13: Using Bézier curves instead of straight lines allows the model to connect the dots in less strokes, and it can draw less jagged shapes, leading to a more aesthetically pleasing result.

### 28 Apr 2026 v2

@@ 1 Introduction
Figure 1: For complex questions, modern chatbots like ChatGPT often return long text responses (a) that are hard for users to understand, verify, and follow. In contrast, SketchVLM guides users (b) step-by-step by annotating the input image and grounding answers to relevant image regions—here, guiding a user how to check their car’s oil level (source: https://www.youtube.com/watch?v=tNNyu9S65E4).
@@ 2 Related Work
Fine-tuned sketching models   MVoT [23] fine-tunes Chameleon [42] to generate interleaved text and image reasoning traces, visualizing intermediate states to support multi-step spatial reasoning. LatentSketchpad [52] and DeepSketcher [51] both move visual reasoning into learned latent or embedding spaces. LatentSketchpad is built with Gemma3 and Qwen2.5-VL-7B, while DeepSketcher uses Qwen2.5-VL-7B. ViLaSR [48] post-trains Qwen-2.5-VL-7B to sketch on the input image with an SVG overlay before responding. Similarly, ThinkMorph is fine-tuned from BAGEL-7B-MoT [15] to generate visual sketches that support its answers, though it directly edits the input image rather than overlaying SVG annotations. Unlike these approaches, we build a harness around SotA VLMs to enable them to annotate on input images by generating a layer of SVGs. Therefore, different from the literature, our approach is training-free and enjoys the generalizability of SotA VLMs to new domains.
@@ 4 Evaluation
7. Ball Drop [Uncaptioned image]   contains 198 synthetically generated images with harder-to-guess answers and ground-truth ball trajectories produced using PHYRE [4]. We generate equal numbers of images with 1, 2, and 3 randomly placed lines, and randomize the ball’s horizontal position. Models must predict the landing container out of four choices and trace the ball trajectory (Secs.˜B.6, 10, D3 and D2).
@@ 5.3 SketchVLM improves counting accuracy [Uncaptioned image]
Experiment   We compare SketchVLMs’ stroke-based annotations against the baseline models that directly output parameters for shape locations and shape size (Sec.˜4.1). To evaluate oval annotations, we convert the rendered shape into its tight enclosing bounding box before computing metrics. Performance is measured using Average Precision (AP) at an IoU threshold of 0.5.
Results   [Uncaptioned image] sketchVLM with stroke-based rectangles is effective for medium and large objects, but remains limited for small-object detection. Sketch-based rectangles slightly improve performance on medium (+0.6) and large objects (+1.4), but significantly degrade small-object detection (-10.1), reducing overall AP50 from 63.1 to 58.8 (Figs.˜6, D10 and C4).
We observe that stroke-based outputs in SketchVLMs underperform the original model on small objects (Tab.˜C4). We examine detection statistics and prompt ablations and find that SketchVLMs match the original model in precision but exhibit lower recall (Tab.˜C5) and do not significantly change with variations in the drawing prompt (Tab.˜C6). Therefore, in order to get the accuracy of pre-defined shapes as well as the expressiveness of free-form strokes, we allow the model to output both simultaneously, such as in Fig.˜1.
@@
Results   SketchVLM improves part labeling for [Uncaptioned image] (+1.3) but slightly underperforms [Uncaptioned image] at strict boundary matching (-3.8) (Tab.˜4). For [Uncaptioned image], SketchVLM becomes increasingly robust under boundary dilation [11], achieving higher accuracy as tolerance increases, with remaining errors dominated by wrong-position mistakes (>79%; Tab.˜C7). For [Uncaptioned image], the gap narrows from (-3.8) at $r=0$ to (-0.6) at $r=7$, reaching near parity under modest tolerance (Tab.˜4). These errors largely correspond to small boundary offsets that are visually negligible (Figs.˜7, 8 and D11), where the original models produce more missing-label errors and SketchVLMs more position errors (Tab.˜C7). These findings indicate that SketchVLMs primarily introduce minor spatial shifts rather than semantic labeling failures, and remain competitive under reasonable boundary tolerance.
@@ 5.6 SketchVLMs outperform models fine-tuned directly on path-tracing tasks [Uncaptioned image]
Results   We find that [Uncaptioned image] and [Uncaptioned image] perform poorly on VPCT with accuracies of 37% and 27%, which are near the random choice baseline of 33.3%. On our Ball Drop dataset, they have accuracies of 35.9% and 30.3%, only slightly above the random choice baseline of 25.0%, showing how these models fail to generalize to tasks that they are not trained on. $\bm{+} +\bm{+} [Uncaptioned image]$ often removes ledges in the image (Fig.˜D1) and performs significantly worse than the baseline [Uncaptioned image] accuracy (Tab.˜2). In contrast, SketchVLMs output coherent annotations and have high accuracy, and [Uncaptioned image] sketchVLM reaches 96.0% accuracy on VPCT and 79.7% on Ball Drop (Fig.˜10).
@@ 5.10 VLM judge ratings positively correlate with human judgments
We validate our annotation quality ratings from our VLM judge (Gemini-3-Flash-Preview [12] (Tab.˜F1)) against three human annotators on 2,250 annotations across Ball Drop, VPCT, and Maze Navigation using a quadratic Kappa score and Pearson Correlation. Agreement between humans and the VLM judge is moderate (quadratic Kappa of $0.51\pm 0.02$, Pearson correlation of $0.52\pm 0.01$) compared to high inter-human agreement (quadratic Kappa of $0.84\pm 0.04$, Pearson correlation of $0.85\pm 0.04$). Further details are in Appendix˜E. While the VLM judge is not as consistent as the human evaluators, its ratings are positively correlated with human judgments across all tasks and models. We therefore treat the VLM judge scores as a useful proxy metric for quick and cost-effective evaluation that should be taken into consideration with both task accuracy and annotation–text alignment.
@@ B.5 Maze Navigation
Given a start point, an end point, and a set of direction commands (e.g., Up, Down, Left, Right), the model must determine if the path reaches the goal without crossing any border walls. We create 200 unique 3x3 grids where the shortest path length between the starting green square and ending red square varies from 3 to 8 steps. For each maze, we take the ground truth path and randomly change one of the direction steps in order to make an invalid path (for example, Left, Right, Down could be changed to Right, Right, Down) (Fig.˜9).
@@ D.6 Part Labeling
Figure D11: Qualitative comparison on part labeling tasks between Refer to caption, Refer to caption sketchVLM , and Refer to caption.
@@ D.8 Connect-the-Dots: Bézier Curves versus Lines
Figure D13: Using Bézier curves instead of straight lines allows the model to connect the dots in fewer strokes, and it can draw less jagged shapes, leading to a more aesthetically pleasing result.